OpenStudy (anonymous):

If (44)x = 432, what is the value of x? 4 7 8 28

OpenStudy (anonymous):

\[4^{4})x = 4^{32}\]

OpenStudy (michele_laino):

is your equation like this? \[\Large {44^x} = 432\]

OpenStudy (anonymous):

no its like the one i just wrote in there

OpenStudy (michele_laino):

sorry!

OpenStudy (anonymous):

its okay (:

OpenStudy (anonymous):

x=8

OpenStudy (anonymous):

your not supposed to give the answer

OpenStudy (michele_laino):

then we have to divide both sides by 4^4, so we get: \[\Large \frac{{{4^4}x}}{{{4^4}}} = \frac{{{4^{32}}}}{{{4^4}}}\] please simplify

OpenStudy (anonymous):

32/ 4 = 8 s therefore x=8

OpenStudy (anonymous):

you have to multiply the exponents that are 4 and x and make them equal to 32 4*x=32 x=32/4 x=8

OpenStudy (michele_laino):

hint: \[\Large \begin{gathered} \frac{{{4^4}}}{{{4^4}}} = {4^4}:{4^4} = {4^{4 - 4}} = {4^0} = ... \hfill \\ \hfill \\ \frac{{{4^{32}}}}{{{4^4}}} = {4^{32}}:{4^4} = {4^{32 - 4}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

michele the answer is x=8 i already got it

OpenStudy (michele_laino):

ok! then your equation is: \[\Large {4^{4x}} = {4^{32}}\]

OpenStudy (anonymous):

so the answer is 8

OpenStudy (michele_laino):

good job! @ana123456

OpenStudy (michele_laino):

correct!